OpenStudy (anonymous):

Solve for x. x^2 + 8x + 16 = 0

OpenStudy (anonymous):

x^2-8x+16 =e^0 x^2 - 8x + 16 = 1 x^2 - 8x + 15 = 0 (x-5) = 0 (x-3) = 0 x = 5 x = 3

OpenStudy (lgbasallote):

break down the middle term x^2 + 4x + 4x + 16 = 0 group (x^2 + 4x) + (4x + 16) = 0 factor out x(x+4) + 4(x+4) = 0 factor again (x+4)(x+4) = 0 solve for x now

OpenStudy (anonymous):

raise both sides to the e power then solve it. (x-5)(x-3)

OpenStudy (lgbasallote):

why is there e^0 lol

OpenStudy (lgbasallote):

what did you do @doubledrive

OpenStudy (lgbasallote):

@derptastic878 do you know what to do from where i stopped?

OpenStudy (anonymous):

lgbasalolote is correct x=-4

OpenStudy (anonymous):

e is for the exponent. im eating and reading the question and using the calc. so typing the e is better than ^. @Igbasallote

OpenStudy (anonymous):

(x+4)(x+4) = (x+4)^2 so sqrt(x+4)^2 = sqrt(0) x+4 = 0 x = -4 Am I right?